Opened 9 years ago

Last modified 3 years ago

#352 accepted flight processing

GB07/09, Flight day 212a/2010, Alps

Reported by: ella Owned by: knpa
Priority: alpha 4 medium Milestone: 2010 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by iopa)

Data location: ~arsf/arsf_data/2010/flight_data/archive-overflow/GB07_09-2010_212a_Alps
Project has been fast copied to ~arsf/workspace.

Data arrived from ARSF via network transfer on 9th August 2010.

Scientific objective: Remote sensing support for research into the processes and rates of supraglacial debris cover formation on alpine glaciers, and on the response of glaciers to increases in supraglacial mass loading.

Priority: a4

PI:Benjamin Brock (and Martin Kirkbridge)
PLEASE NOTE:This project has a dual PI and we should cc Matin Kirkbridge on all emails. Data delivery should also be offered to Martin Kirkbridge, too.

Sensors:

  • Eagle (12/05/11)
  • Hawk (12/05/11)
  • Leica LIDAR (20/09/10)
  • RCD Camera (16/09/10)

Attachments (1)

2122010.sync (207 bytes) - added by iopa 9 years ago.
Eagle & Hawk SCT offsets

Download all attachments as: .zip

Change History (44)

comment:1 Changed 9 years ago by ella

  • Description modified (diff)

comment:2 Changed 9 years ago by ella

  • Description modified (diff)

comment:3 Changed 9 years ago by ella

  • Description modified (diff)

comment:4 Changed 9 years ago by jb362

Starting hyperspectral processing.

comment:5 Changed 9 years ago by jb362

I'm leaving the hyperspectral processing until we have processed the LiDAR for this project since the SRTM dem is useless in such steep terrain.

comment:6 Changed 9 years ago by ella

IPAS navigation now processed.
[Applanix already completed by jb362]

comment:7 Changed 9 years ago by emca

  • Owner set to emca
  • Status changed from new to accepted

Starting LiDAR processing

comment:8 Changed 9 years ago by iopa

Starting camera check.

comment:9 Changed 9 years ago by iopa

Camera check:

1) delete camera delivery log from photographs dir
2) exiftool sample run has used the camera_delivery.log in the readme, readme needs re-creating

All else seems correct.

Note: readme front page contains both the name of the PI as given in the application and the name of collaborator 2 who was entered as the PI in the logsheet.

comment:10 Changed 9 years ago by jb362

Camera ready for delivery.

comment:11 Changed 9 years ago by jb362

Iopa will be delivering this for me since I'm leaving.

comment:12 Changed 9 years ago by emca

LiDAR data ready for delivery check at
~airborne/workspace/GB07_09-2010_212a_Alps/delivery/20100910/GB07_09

comment:13 Changed 9 years ago by iopa

  • Description modified (diff)

RCD camera was dispatched today, 16th September 2010.

comment:14 Changed 9 years ago by iopa

Delivery checking LiDAR

comment:15 Changed 9 years ago by iopa

Delivery check report;

All is good and ready to deliver.

comment:16 Changed 9 years ago by emca

  • Description modified (diff)

LiDAR data sent 20/09/2010

comment:17 Changed 9 years ago by emca

  • Status changed from accepted to assigned

LiDAR and Camera data have been rsync'd to
~arsf/arsf_data/2010/flight_data/western_europe/GB07_09-2010_212a_Alps

comment:18 Changed 9 years ago by knpa

Appears to be an issue with the bandset labels in the header files; they start at 900nm which is lower than it should be. Possibly incorrect bandset selected or none specified. Will leave this until end of the year and new Hawk calibration.

comment:19 Changed 9 years ago by benj

  • Description modified (diff)

comment:20 Changed 9 years ago by iopa

Taking hyperspectral.

comment:21 Changed 9 years ago by iopa

Hyperspectral status:

All data has been run thhrough azgcorr but the level 3's show tearings and various bad effects caused by the low resolution of the SRTM DEM that was then patched with the LiDAR
and the patching introduced great discontinuities. SRTM-only DEM previously showed even
worse effects.

In progress was the use of an ASTER DEM which would be patched to LiDAR, but before that it would be interpolated with 'cubic convolution' (through ENVI which can do this for BIL DEMs, GRASS can output BIL and then ENVI can do the interpolation and then back to GRASS for the patching),but while the interpolation succeeded the patching myself I did not succeed with and that is where this stands.

comment:22 Changed 9 years ago by iopa

Hyperspectral has been processed and delivery is ready for checking.

@ /users/rsg/arsf/workspace/GB07_09-2010_212a_Alps/delivery/20110511/GB07_09

comment:23 Changed 9 years ago by emca

Taking hyperspectral delivery check

comment:24 Changed 9 years ago by emca

Eagle/Hawk line 1 does not process without -bend, can you remove the offending lines from the lev1 files

You've included the datum/correction file in the delivery, but its not used in the sample command - is this file needed to process the data?

Change the project code on the logsheet to GB07/09

Dem name needs to be *-aster.dem to be comply to standard

Add the SCT's to the ticket

Otherwise data looks good

Changed 9 years ago by iopa

Eagle & Hawk SCT offsets

comment:25 Changed 9 years ago by iopa

Delivery fixes have been made, data is getting ready for dispatch.

comment:26 Changed 9 years ago by iopa

Hyperspectral dispatched to PI Ben Brock on 12th May 2011.

comment:27 Changed 9 years ago by iopa

  • Description modified (diff)

Project has been tidied and r-synced to the repository.

comment:28 Changed 9 years ago by emca

  • Component changed from Processing: general to Archiving

comment:29 Changed 8 years ago by knpa

Reprocessing hyperspectral with APL.

comment:30 Changed 8 years ago by knpa

Stages 1&2 completed.

comment:31 Changed 8 years ago by knpa

Stage 3 tested successfully.

There is an issue with hawk line 7 however that stopped this moving beyond lev1 - apl claims no matching sync message in nav file that matches hdr times.

comment:32 Changed 8 years ago by knpa

Fixed this issue.

comment:33 Changed 8 years ago by knpa

Deleted delivery and reprocessing as apl chain has been updated to generate additional masked files.

comment:34 Changed 8 years ago by knpa

Re-ran stage 2 for Hawk as aplcal has been updated in relation to the bad pixel file.

comment:35 Changed 7 years ago by knpa

ReadMe created, ready for delivery check.

comment:36 Changed 7 years ago by emca

Delivery checked reprocessed apl data. Lev3 data looks a bit mismatched in some areas (where there is no lidar overlap) but this is due lack of good DEM coverage.

  • Flight letter missing from Readme (project details + delivery contents)
  • Readme: Level 3 (pg 4) gdal_translate sample command and paragraph after sample command problem with band list ("-b b -b a -b n", "b, a, n, d, s, , 3, 0, „ , 1, 5, „ , 7")
  • DEM and logsheet named incorrectly (- vs. _)

comment:37 Changed 7 years ago by knpa

  • Owner changed from emca to knpa
  • Status changed from assigned to accepted

comment:38 Changed 7 years ago by knpa

  • Milestone set to 2010 data processing completion

Fixed above points.

Delivery of reprocessed HS has not been requested so this is finished.

comment:39 Changed 6 years ago by knpa

Merged and deleted workspace version.

comment:40 Changed 4 years ago by dac

Archiving

Started uploading camera, LiDAR and raw data to NEODC. Reprocessed hyperspectral already archived.

comment:41 Changed 4 years ago by dac

Archiving

Data uploaded to NEODC.

comment:42 Changed 4 years ago by dac

Archiving

Processed data available from NEODC (http://browse.ceda.ac.uk/browse/neodc/arsf/2010/GB07_09/GB07_09-2010_212a_Alps) - raw data uploaded but not yet archived.

comment:43 Changed 3 years ago by dac

Archiving

Raw data now available from NEODC.

Note: See TracTickets for help on using tickets.